# SecAlign: Defending Against Prompt Injection with Preference Optimization

### v2

@@ Abstract.
To mitigate this vulnerability, we propose a new defense called SecAlign based on the technique of preference optimization. Our defense first constructs a preference dataset with prompt-injected inputs, secure outputs (ones that respond to the legitimate instruction), and insecure outputs (ones that respond to the injection). We then perform preference optimization on this dataset to teach the LLM to prefer the secure output over the insecure one. This provides the first known method that reduces the success rates of various prompt injections to around 0%, even against attacks much more sophisticated than ones seen during training. This indicates our defense generalizes well against unknown and yet-to-come attacks. Also, our defended models are still practical with similar utility to the one before our defensive training. Our code is here.
@@
copyright: acmlicensedjournalyear: 2025ccs: Security and privacy Systems security
@@ 1. Introduction
Figure 1. Top: We formulate defense against prompt injection as a preference optimization problem. Given a prompt-injected input with the injected instruction highlighted in red, the LLM is fine-tuned to prefer the response to the instruction over the response to the injection. Bottom: Our proposed SecAlign reduces the attack success rate of the strongest tested prompt injection to around 0% without hurting the utility from Llama3-8B-Instruct [Dubey et al., 2024], an advanced LLM. In comparison, state-of-the-art (SOTA) prompting-based defense In-Context [Wei et al., 2024], see Table 2, and fine-tuning-based defense StruQ [Chen et al., 2025] achieve very limited security with utility loss.
@@
Intuitively, prompt injection attacks exploit the inability of LLMs to distinguish between instruction (from a trusted system designer) and data (from an untrusted user) in their input. Existing defenses try to explicitly enforce the separation between instruction and data via prompting [202, 2023a, Willison, 2023a, Liu et al., 2024] or fine-tuning [Yi et al., 2023, Piet et al., 2023, Chen et al., 2025, Wallace et al., 2024, Wu et al., 2025]. Fine-tuning defenses, which are empirically validated to be stronger in prior work [Chen et al., 2025], adopt a training loss that maximizes LLM’s likelihood of outputting the desirable response (to the benign instruction) under prompt injection, so that the injected instruction is ignored.
@@
Unfortunately, existing defenses are brittle against attacks that are unseen in fine-tuning time. For example, StruQ [Chen et al., 2025] suffers from over 50% attack success rate under an attack that optimizes the injection [Zou et al., 2023]. This lack of generalization against unseen attacks makes existing defenses fragile, since attackers are motivated to continue evolving their techniques. We show that the fragility of existing fine-tuning-based defenses may stem from an underspecification in the fine-tuning objective: The LLM is only trained to favor the desirable response, but does not know what an undesirable response looks like. Thus, a secure LLM should also observe the response to the injected instruction and be steered away from that response. Coincidentally, this learning problem is well-studied under the name of preference optimization, and is commonly used to align LLMs to human preferences such as ethics and discrimination
@@
This leads us to formulate prompt injection defense as preference optimization: given a prompt-injected input $x$, the LLM is fine-tuned to prefer the response $y_{w}$ to the instruction over the response $y_{l}$ to the injection; see Fig. 1 (top). We then propose our method, called SecAlign, which builds a preference dataset with input-desirable_response-undesirable_response $\{(x,y_{w},y_{l})\}$ triples, and performs preference optimization on it. Similar to the idea of using preference optimization for aligning to human values, we demonstrate that ”security against prompt injection” is also a preference that could be optimized, which, interestingly, requires no human labor vs. alignment due to the well-defined prompt injection security policy.
@@
We evaluate SecAlign against three (strongest out of a dozon ones tested in [Chen et al., 2025]) optimization-free prompt injection attacks and three optimization-based attacks (GCG [Zou et al., 2023], AdvPrompter [Paulus et al., 2024], and NeuralExec [Pasquini et al., 2024]) on seven models. SecAlign maintains the same level of utility as the non-preference-optimized counterpart no matter whether the preference dataset is in a same or different domain as instruction tuning. More importantly, SecAlign achieves SOTA security with consistent 0% optimization-free attack success rates (ASRs). For stronger optimization-based attacks, SecAlign achieves the ASR mainly around 0% for the first time to our knowledge, and consistently reduces the ASR by a factor of ¿4 from the current SOTA StruQ [Chen et al., 2025]. In comparison, see Fig. 1 (bottom), existing SOTA prompting-based or fine-tuning-based defenses have limited security with optimization-based ASRs consistently over 40%.
@@ 2.2. Problem Significance
In general, prompt injection attacks can lead to leakage of sensitive information and privacy breaches, and will likely severely limit deployment of LLM-integrated applications if left unchecked, which has also been shown in other productions such as Google Bard [202, 2023b], Anthropic Web Agent [202, 2024a], and OpenAI ChatGPT [202, 2024b]. To enable new opportunities for safely using LLMs in systems, our goal is to design fundamental defenses that are robust to advanced LLM prompt injection techniques. A comprehensive solution has not yet been developed. Among recent progress [Liu et al., 2024, Yi et al., 2023, Suo, 2024, Rai et al., 2024, Yip et al., 2023, Piet et al., 2023], Piet et al. [2023], Chen et al. [2025] show promising robustness against optimization-free prompt injections, but none of them are robust to optimization-based prompt injections. Recently, Wallace et al. [2024] introduces the instruction hierarchy, a generalization of [Chen et al., 2025], which aims to always prioritize the instruction with a high priority if it conflicts with the low-priority instruction, e.g., injected prompt in the data. OpenAI deployed the instruction hierarchy [Wallace et al., 2024] in GPT-4o mini, a frontier LLM. It does not use any undesirable samples to defend against prompt injections like SecAlign, despite their usage of alignment training to consider human preferences.
@@ 2.3. Optimization-Free Prompt Injections
We first introduce manually-designed prompt injections, which have a fixed format with a clear attack intention. We denote them as optimization-free as these attacks are constructed manually rather than through iterative optimization. Among over a dozen optimization-free prompt injections introduced in [Chen et al., 2025], the below ones are the strongest or most representative, so we use them in our method design (training) or evaluation (testing). Among all described attacks in this section, we only train the model with simple Naive and Completion attacks, but test it with all attacks to evaluate model’s defense performance on unknown sophisticated attacks, especially on strong optimization-based ones.
Naive Attack.
Naive attack directly puts the injected prompt inside the data [Liu et al., 2024].
A prompt injection example by Naive attack $d_{\text{instruction}}$ Please generate a python function for the provided task. $d_{\text{data}}$ Determine whether a number is prime. Do dinosaurs exist? $d_{\text{response}}$
@@
Generally, the attacker wants to highlight the injected prompt to the LLM, and asks explicitly the LLM to follow this new instruction. This leads to an Ignore attack [Perez and Ribeiro, 2022], which includes some deviation sentences (e.g., “Ignore previous instructions and …”) before the injected prompt. An example is in Section 2.1. We randomly choose one of the ten deviation sentences designed in [Chen et al., 2025] to attack each sample in our evaluation.
@@
Completion attacks are very effective [Chen et al., 2025, Liu et al., 2024]. We can also combine Ignore and Combination attacks to perform a Ignore-Completion attack.
@@ 3.1. Revisiting Prompt Injection Defenses
On the other side, there are two complementary objectives for prompt injection defense: (i) encouraging the desirable output by fine-tuning the LLM to maximize the likelihood of $y_{w}$; and (ii) discouraging the undesirable output by minimizing the likelihood of $y_{l}$. Existing defenses [Yi et al., 2023, Chen et al., 2025, Wallace et al., 2024, Wu et al., 2025] only aim for (i) following adversarial training (AT) [Madry et al., 2018], by far the most effective defense for classifiers, to mitigate prompt injection. That is, minimize the standard training loss on attacked (prompt-injected) samples $x$:
@@
Targeting only at (i) when securing LLMs as in securing classifiers neglects the difference between these two types of models. For classifiers, encouraging prediction on $y^{*}$ is almost equivalent to discouraging prediction on $y^{\prime}$ because the number of possible predictions is small. For LLMs, however, objectives (i) and (ii) are only loosely correlated: An LLM typically has a vocabulary size $V$ and an output length $L$, leading to $V^{L}$ possible outputs. Due to the exponentially larger space of LLM outputs, regressing an LLM towards a $y_{w}$ has limited influence on LLM’s probability to output a large number of other sentences, including $y_{l}$. This explains why existing fine-tuning-based defenses [Chen et al., 2025, Yi et al., 2023, Wallace et al., 2024, Wu et al., 2025] suffer from over $50\%$ attack success rates: the loss Eq. 1 only specifies objective (i), which cannot lead to the achievement of (ii) in fine-tuning LLMs.
@@ 3.3. Implementing SecAlign: Preference Dataset
To craft SecAlign preference dataset, we need to format the instruction and data $s$ into one input string for LLMs, see also Section 2.1. To enforce security under prompt injections in an AT-style, the input should be attacked (prompt-injected), so we put an instruction at the end of the data part following [Chen et al., 2025]. The injected instruction comes from another random sample (e.g., $s^{\prime}$) in the instruction tuning dataset, so we do not need to manually write injections as in [Yi et al., 2023]. For the output, the security policy of prompt injections asks the LLM to respond to the benign instruction instead of the injected instruction. Thus, the ”desirable output” is the response to the benign instruction in $s$. The ”undesirable output” is the response to the injected instruction, which, interestingly, turns out to be the ”desirable output” in $s^{\prime}$ where the injection is from.
@@
We summarize our procedure to construct the preference dataset in Algorithm 1 with more details. In our implementation, we mostly (90%) prompt-inject the input by the Naive attack as the above examples, but additionally do Completion attacks (10%) to get better defense performance as recommended by [Chen et al., 2025], which also offers us hundreds of additional delimiters ($d^{\prime}_{\text{instruction}}$, $d^{\prime}_{\text{data}}$, $d^{\prime}_{\text{response}}$) to diversify the Completion attack. As in Section 2.3, a Completion attack manipulates the input structure by adding delimiters $d^{\prime}$ to mimic the conversation, see Lines 8-10 in Algorithm 1.
@@
6:        $s_{\text{data}}$ += $s^{\prime}_{\text{instruction}}+s^{\prime}_{\text{data}}$        # Naive attack
@@
8:        Sample attack delimiters $d^{\prime}$ from [Chen et al., 2025]   # Completion attack
@@
(4) $\min_{\theta}\mathop{\mathbb{E}}_{(\hat{x},y)}\mathcal{L}(\theta,x,y)=\min_{\theta}\mathop{\mathbb{E}}_{(\hat{x},y)}\left[\max_{x\in\mathcal{C}}\mathcal{L}(\theta,x,y)\right],$
where $x$ represents the attacked example constructed from the original sample $\hat{x}$ by solving the inner optimization (under constraint $C$) to simulate an attack. Let us re-write Eq. 3 as
@@
Instead of optimizing the attacked sample $x$ by gradients
$\min_{\theta}\mathop{\mathbb{E}}_{(\hat{x},y)}\left[\max_{x\in\mathcal{C}}\mathcal{L}_{\text{SecAlign}}(\theta,x,y)\right]$
as in Eq. 4, SecAlign resorts to optimization-free attacked samples $x$ to loosely represent the maximum.
(5) $\min_{\theta}\mathop{\mathbb{E}}_{(x,y)}\mathcal{L}(\theta,x,y).$
This is because existing optimizers for LLMs like GCG [Zou et al., 2023] cannot work within a reasonable time budget (hundreds of GPU hours) for training. Besides, optimization-free attacks like Completion attacks have been shown effective in prompt injections [Chen et al., 2025] and could be an alternative way to maximize the training loss.
@@ 4.1. Experimental Setup
Following [Chen et al., 2025], we use the whole AlpacaFarm dataset [Dubois et al., 2024] to evaluate utility, and its samples with a data part (when prompt injection applies) to evaluate security. AlpacaFarm is an instruction tuning dataset [Dubois et al., 2024] with 805 well-designed general-purpose samples, among which 208 have a data part. We use the Cleaned Alpaca instruction tuning dataset [Ruebsamen, 2024] to generate the preference dataset for training. AlpacaFarm [Dubois et al., 2024] is in another domain as Cleaned Alpaca dataset [Ruebsamen, 2024]. Despite having similar names, they are essentially two datasets instead of splits from one dataset, and their samples are without overlap inherently. The community is thus using AlpacaFarm to evaluate LLMs trained on Cleaned Alpaca, see [Li et al., 2023a, Chen et al., 2025, Wu et al., 2025].
@@
We use AlpacaEval2 [Li et al., 2023a], an LLM-as-a-judge-based evaluation of a model’s general-purpose utility, to assess our model. It runs the model on all 805 AlpacaFarm samples, gets reference responses from a reference model (davinci003 in our experiments), and uses GPT-4-turbo to compare the reference responses with those from the LLM-under-test with a specially-crafted prompt. A WinRate of 100% indicates that the LLM-under-test is consistently better than the reference model, and 50% means the two LLMs are equivalent in output quality. Note that AlpacaEval2 ranking is very close to human evaluations, enjoying a 0.98 correlation to Chatbot Arena [Chiang et al., 2024], a benchmark using human feedback.
@@
We use the attack success rate of prompt injections to quantify the security. We evaluate the attack success rate of all attacks in Section 2. We report the Maximum Attack Success Rate (Max ASR) of all tested optimization-free injection attacks (Ignore, Completion, Ignore-Completion) and that of all tested optimization-based injection attacks (GCG, AdvPrompter, and NeuralExec), with breakdown numbers in the Appendix. All test attacks have never been seen in training, which only uses simple Naive and Completion with a different attack delimiter set as in test-time Completion attacks following [Chen et al., 2025]. We use each of those prompt injections to attack 208 AlpacaFarm [Dubois et al., 2024] samples that have a data item. AlpacaFarm favors attackers as it contains shorter data than those in prior studies [Liu et al., 2024, Yi et al., 2023], making the injection more prominent. Following [Chen et al., 2025], the evaluation prompt injection, if not otherwise stated, is “Print exactly Hacked!”, which has never appeared in the simulated injections seen in training, which is sampled inside the training instruction tuning dataset [Ruebsamen, 2024]. We inject “Print exactly Hacked!” at the end of the data, which favors attackers more than any other location, see Table 4 and [Piet et al., 2023, Yi et al., 2023]. The attack is regarded as successful if and only if the response begins with “Hacked” or “hacked” if not otherwise stated.
@@
We apply SecAlign to seven open-weight models [Inc, 2021]: Mistral-7B-Instruct [Jiang et al., 2023], Llama3-8B-Instruct [Dubey et al., 2024], Llama-7B [Touvron et al., 2023b], Mistral-7B [Jiang et al., 2023], Llama3-8B [Dubey et al., 2024], Yi-1.5-6B [Young et al., 2024], and Llama-13B [Touvron et al., 2023b]. The first two models have been SFT-ed with their private commercial instruction tuning datasets, so we could directly apply SecAlign on them with their offered delimiters: system delimiter for the instruction, user delimiter for the data, and assistant delimiter for the response. The last five are base pretrained models and should be SFTed before DPO [Rafailov et al., 2024], so we perform standard (non-defensive) SFT following [Chen et al., 2025], which reserves three special tokens for each of the delimiters. That is, $d_{\text{instruction}}=$[MARK] [INST] [COLN], where each token above has a unique trainable embedding vector during the model tokenization, and similarly for $d_{\text{data}}$ and $d_{\text{response}}$.
@@
In DPO, we use sigmoid activation $\sigma$ and $\beta=0.1$ as the default recommendation. Due to the involvement of two checkpoints $\pi_{\theta},\pi_{\mathrm{ref}}$ in DPO Eq. 3, the memory consumption almost doubles. To ease the training, we adopt LoRA [Hu et al., 2022], a memory efficient fine-tuning technique that only optimizes a very small proportion ($<0.5\%$ in all our studies) of the weights but enjoys performance comparable to fine-tuning the whole model. The LoRA hyperparameters are r=64, lora_alpha=8, lora_dropout=0.1, target_modules = ["q_proj", "v_proj"]. We use the TRL library [von Werra et al., 2020] to implement DPO, and Peft library [Mangrulkar et al., 2022] to implement LoRA. Our training requires 4 NVIDIA Tesla A100s (80GB) to support Pytorch FSDP [Zhao et al., 2023]. We perform DPO for 3 epochs with the tuned learning rates $[1.4,1.6,2,1.4,1.6,1.6,2]\times 10^{-4}$ for the seven models above respectively. In standard SFT (required before SecAlign for base models) and defensive SFT (the precise StruQ defense [Chen et al., 2025]), we fine-tune the LLMs for 3 epochs using the learning rate $[20,2.5,2,2,20]\times 10^{-6}$ for the five base models above respectively.

### v3

@@ Abstract.
To mitigate this vulnerability, we propose a new defense called SecAlign based on the technique of preference optimization. Our defense first constructs a preference dataset with prompt-injected inputs, secure outputs (ones that respond to the legitimate instruction), and insecure outputs (ones that respond to the injection). We then perform preference optimization on this dataset to teach the LLM to prefer the secure output over the insecure one. This provides the first known method that reduces the success rates of various prompt injections to ¡10%, even against attacks much more sophisticated than ones seen during training. This indicates our defense generalizes well against unknown and yet-to-come attacks. Also, SecAlign models are still practical with similar utility to the one before defensive training in our evaluations. Our code is here.
@@
journalyear: 2025copyright: rightsretainedconference: Proceedings of the 2025 ACM SIGSAC Conference on Computer and Communications Security; October 13–17, 2025; Taipei, Taiwan.booktitle: Proceedings of the 2025 ACM SIGSAC Conference on Computer and Communications Security (CCS ’25), October 13–17, 2025, Taipei, Taiwanisbn: 979-8-4007-1525-9/2025/10doi: 10.1145/3719027.3744836copyright: acmlicensedjournalyear: 2025ccs: Security and privacy Systems security
@@ 1. Introduction
Figure 1. Top: We formulate defense against prompt injection as a preference optimization problem. Given a prompt-injected input with the injected instruction highlighted in red, the LLM is fine-tuned to prefer the response to the instruction over the response to the injection. Bottom: Our proposed SecAlign reduces the attack success rate of the strongest tested prompt injection to 8% without hurting the utility from Llama3-8B-Instruct [Dubey et al., 2024], an advanced LLM. In comparison, state-of-the-art (SOTA) prompting-based defense In-Context [Wei et al., 2024], see Table 2, and fine-tuning-based defense StruQ [Chen et al., 2025a] achieve very limited security with utility loss.
@@
Intuitively, prompt injection attacks exploit the inability of LLMs to distinguish between instruction (from a trusted system designer) and data (from an untrusted user) in their input. Existing defenses try to explicitly enforce the separation between instruction and data via prompting [202, 2023a, Willison, 2023a, Liu et al., 2024] or fine-tuning [Yi et al., 2023, Piet et al., 2023, Chen et al., 2025a, Wallace et al., 2024, Wu et al., 2025a]. Fine-tuning defenses, which are empirically validated to be stronger in prior work [Chen et al., 2025a], adopt a training loss that maximizes LLM’s likelihood of outputting the desirable response (to the benign instruction) under prompt injection, so that the injected instruction is ignored.
@@
Unfortunately, existing defenses are brittle against attacks that are unseen in fine-tuning time. For example, StruQ [Chen et al., 2025a] suffers from over 50% attack success rate under an attack that optimizes the injection [Zou et al., 2023]. This lack of generalization against unseen attacks makes existing defenses fragile, since attackers are motivated to continue evolving their techniques. We show that the fragility of existing fine-tuning-based defenses may stem from an underspecification in the fine-tuning objective: The LLM is only trained to favor the desirable response, but does not know what an undesirable response looks like. Thus, a secure LLM should also observe the response to the injected instruction and be steered away from that response. Coincidentally, this learning problem is well-studied under the name of preference optimization, and is commonly used to align LLMs to human preferences such as ethics and discrimination.
@@
This leads us to formulate prompt injection defense as preference optimization: given a prompt-injected input $x$, the LLM is fine-tuned to prefer the response $y_{w}$ to the instruction over the response $y_{l}$ to the injection; see Fig. 1 (top). We then propose our method, called SecAlign, which builds a preference dataset with input-desirable_response-undesirable_response $\{(x,y_{w},y_{l})\}$ triples, and performs preference optimization on it. Similar to the idea of using preference optimization for aligning to human values, we demonstrate that ”security against prompt injection” is also a preference that could be optimized, which, interestingly, requires no human labor vs. alignment (to human preference) due to the well-defined prompt injection security policy.
@@
We evaluate SecAlign against three (strongest ones out of a dozon ones tested in [Chen et al., 2025a]) optimization-free prompt injection attacks and three optimization-based attacks (GCG [Zou et al., 2023], AdvPrompter [Paulus et al., 2024], and NeuralExec [Pasquini et al., 2024]) on five models. SecAlign maintains the same level of utility as the non-preference-optimized counterpart no matter whether the preference dataset is in a same or different domain as instruction tuning. More importantly, SecAlign achieves SOTA security with consistent 0% optimization-free attack success rates (ASRs). For stronger optimization-based attacks, SecAlign achieves the ASR mainly ¡10% for the first time to our knowledge, and consistently reduces the ASR by a factor of ¿4 from the current SOTA StruQ [Chen et al., 2025a]. In comparison, see Fig. 1 (bottom), existing SOTA prompting-based or fine-tuning-based defenses have limited security with optimization-based ASRs consistently over 40%.
Following this work, we use an improved SecAlign to build the first open-source commercial-grade (70B) LLM with built-in defense against prompt injection attacks [Chen et al., 2025b], which is more robust than existing industry solutions especially in agentic settings where prompt injection security is a priority.
@@ 2.2. Problem Significance
In general, prompt injection attacks can lead to leakage of sensitive information and privacy breaches, and will likely severely limit deployment of LLM-integrated applications if left unchecked, which has also been shown in other productions such as Google Bard [202, 2023b], Anthropic Web Agent [202, 2024a], and OpenAI ChatGPT [202, 2024b]. To enable new opportunities for safely using LLMs in systems, our goal is to design fundamental defenses that are robust to advanced LLM prompt injection techniques. A comprehensive solution has not yet been developed. Among recent progress [Liu et al., 2024, Yi et al., 2023, Suo, 2024, Rai et al., 2024, Yip et al., 2023, Piet et al., 2023], Piet et al. [2023], Chen et al. [2025a] show promising robustness against optimization-free prompt injections, but none of them are robust to optimization-based prompt injections. Recently, Wallace et al. [2024] introduces the instruction hierarchy, a generalization of [Chen et al., 2025a], which aims to always prioritize the instruction with a high priority if it conflicts with the low-priority instruction, e.g., injected prompt in the data. OpenAI deployed the instruction hierarchy [Wallace et al., 2024] in GPT-4o mini, a frontier LLM. It does not use any undesirable samples to defend against prompt injections like SecAlign, despite their usage of alignment training to consider human preferences.
@@ 2.3. Optimization-Free Prompt Injections
We first introduce manually-designed prompt injections, which have a fixed format with a clear attack intention. We denote them as optimization-free as these attacks are constructed manually rather than through iterative optimization. Among over a dozen optimization-free prompt injections introduced in [Chen et al., 2025a], the below ones are the strongest or most representative, so we use them in our method design (training) or evaluation (testing). Among all described attacks in this section, we only train the model with simple Straightforward and Completion attacks, but test it with all attacks to evaluate model’s defense performance on unknown sophisticated attacks, especially on strong optimization-based ones.
Straightforward Attack.
Straightforward attack directly puts the injected prompt inside the data [Liu et al., 2024].
A prompt injection example by Straightforward attack $d_{\text{instruction}}$ Please generate a python function for the provided task. $d_{\text{data}}$ Determine whether a number is prime. Do dinosaurs exist? $d_{\text{response}}$
@@
Generally, the attacker wants to highlight the injected prompt to the LLM, and asks explicitly the LLM to follow this new instruction. This leads to an Ignore attack [Perez and Ribeiro, 2022], which includes some deviation sentences (e.g., “Ignore previous instructions and …”) before the injected prompt. An example is in Section 2.1. We randomly choose one of the ten deviation sentences designed in [Chen et al., 2025a] to attack each sample in our evaluation.
@@
Completion attacks are very effective [Chen et al., 2025a, Liu et al., 2024]. We can also combine Ignore and Combination attacks to perform a Ignore-Completion attack.
@@ 3.1. Revisiting Prompt Injection Defenses
On the other side, there are two complementary objectives for prompt injection defense: (i) encouraging the desirable output by fine-tuning the LLM to maximize the likelihood of $y_{w}$; and (ii) discouraging the undesirable output by minimizing the likelihood of $y_{l}$. Existing defenses [Yi et al., 2023, Chen et al., 2025a, Wallace et al., 2024, Wu et al., 2025a] only aim for (i) following adversarial training (AT) [Madry et al., 2018], by far the most effective defense for classifiers, to mitigate prompt injection. That is, minimize the standard training loss on attacked (prompt-injected) samples $x$:
@@
Targeting only at (i) when securing LLMs as in securing classifiers neglects the difference between these two types of models. For classifiers, encouraging prediction on $y^{*}$ is almost equivalent to discouraging prediction on $y^{\prime}$ because the number of possible predictions is small. For LLMs, however, objectives (i) and (ii) are only loosely correlated: An LLM typically has a vocabulary size $V$ and an output length $L$, leading to $V^{L}$ possible outputs. Due to the exponentially larger space of LLM outputs, regressing an LLM towards a $y_{w}$ has limited influence on LLM’s probability to output a large number of other sentences, including $y_{l}$. This explains why existing fine-tuning-based defenses [Chen et al., 2025a, Yi et al., 2023, Wallace et al., 2024, Wu et al., 2025a] suffer from over $50\%$ attack success rates: the loss Eq. 1 only specifies objective (i), which cannot lead to the achievement of (ii) in fine-tuning LLMs.
@@ 3.3. Implementing SecAlign: Preference Dataset
To craft SecAlign preference dataset, we need to format the instruction and data $s$ into one input string for LLMs, see also Section 2.1. To enforce security under prompt injections in an AT-style, the input should be attacked (prompt-injected), so we put an instruction at the end of the data part following [Chen et al., 2025a]. The injected instruction comes from another random sample (e.g., $s^{\prime}$) in the instruction tuning dataset, so we do not need to manually write injections as in [Yi et al., 2023]. For the output, the security policy of prompt injections asks the LLM to respond to the benign instruction instead of the injected instruction. Thus, the ”desirable output” is the response to the benign instruction in $s$. The ”undesirable output” is the response to the injected instruction, which, interestingly, turns out to be the ”desirable output” in $s^{\prime}$ where the injection is from.
@@
We summarize our procedure to construct the preference dataset in Algorithm 1 with more details. In our implementation, we mostly (90%) prompt-inject the input by the Straightforward attack as the above examples, but additionally do Completion attacks (10%) to get better defense performance as recommended by [Chen et al., 2025a], which also offers us hundreds of additional delimiters ($d^{\prime}_{\text{instruction}}$, $d^{\prime}_{\text{data}}$, $d^{\prime}_{\text{response}}$) to diversify the Completion attack. As in Section 2.3, a Completion attack manipulates the input structure by adding delimiters $d^{\prime}$ to mimic the conversation, see Lines 8-10 in Algorithm 1.
@@
6:        $s_{\text{data}}$ += $s^{\prime}_{\text{instruction}}+s^{\prime}_{\text{data}}$        # Straightforward attack
@@
8:        Sample attack delimiters $d^{\prime}$ from [Chen et al., 2025a]   # Completion attack
@@
(4) $\min_{\theta}\mathop{\mathbb{E}}_{(\hat{x},y)}\left[\max_{x\in\mathcal{C}(\hat{x})}\mathcal{L}(\theta,x,y)\right],$
where $x$ represents the attacked example constructed from the original sample $\hat{x}$ by solving the inner optimization (under constraint $\mathcal{C}$) to simulate an attack. Let us re-write Eq. 3 as
@@
Instead of optimizing the attacked sample $x$ by gradients as in Eq. 4, SecAlign resorts to optimization-free attack $\mathcal{A}$ on the original sample $\hat{x}$ to loosely represent the inner maximum.
(5) $\min_{\theta}\mathop{\mathbb{E}}_{(\hat{x},y)}\mathcal{L}_{\text{SecAlign}}(\theta,\mathcal{A}(\hat{x}),y).$
This is because existing optimizers for LLMs like GCG [Zou et al., 2023] cannot work within a reasonable time budget (hundreds of GPU hours) for training. Besides, optimization-free attacks like Completion attacks have been shown effective in prompt injections [Chen et al., 2025a] and could be an alternative way to maximize the training loss.
@@ 4.1. Experimental Setup
Following [Chen et al., 2025a], we use the whole AlpacaFarm dataset [Dubois et al., 2024] to evaluate utility, and its samples with a data part (when prompt injection applies) to evaluate security. AlpacaFarm is an instruction tuning dataset [Dubois et al., 2024] with 805 well-designed general-purpose samples, among which 208 have a data part. We use the Cleaned Alpaca instruction tuning dataset [Ruebsamen, 2024] to generate the preference dataset for training. AlpacaFarm [Dubois et al., 2024] is in another domain as Cleaned Alpaca dataset [Ruebsamen, 2024]. Despite having similar names, they are essentially two datasets instead of splits from one dataset, and their samples are without overlap inherently. The community is thus using AlpacaFarm to evaluate LLMs trained on Alpaca [Li et al., 2023a, Chen et al., 2025a, Wu et al., 2025a].
@@
We use AlpacaEval2 [Li et al., 2023a], an LLM-as-a-judge-based evaluation of a model’s general-purpose utility, to assess our model. It runs the model on all 805 AlpacaFarm samples, gets reference responses from a reference model (davinci003 in our experiments), and uses GPT-4-turbo to compare the reference responses with those from the LLM-under-test with a specially-crafted prompt. A WinRate of 100% indicates that the LLM-under-test is consistently better than the reference model, and 50% means the two LLMs are equivalent in output quality.
@@
We use the attack success rate of prompt injections to quantify the security. We evaluate the attack success rate of all attacks in Section 2. We report the Maximum Attack Success Rate (Max ASR) of all tested optimization-free injection attacks (Ignore, Completion, Ignore-Completion) and that of all tested optimization-based injection attacks (GCG, AdvPrompter, and NeuralExec), with breakdown numbers in the Appendix. All test attacks have never been seen in training, which only uses simple Straightforward and Completion with a different attack delimiter set as in test-time Completion attacks following [Chen et al., 2025a]. All optimization-based attacks against SecAlign are performed on the LLM after defensive fine-tuning, and with the test-time prompt template.
We use each of those prompt injections to attack 208 AlpacaFarm [Dubois et al., 2024] samples that have a data item. AlpacaFarm favors attackers as it contains shorter data than those in prior studies [Liu et al., 2024, Yi et al., 2023], making the injection more prominent. Following [Chen et al., 2025a], the evaluation prompt injection, if not otherwise stated, is “Print exactly Hacked!”, which has never appeared in the simulated injections seen in training, which is sampled inside the training instruction tuning dataset [Ruebsamen, 2024]. We inject “Print exactly Hacked!” at the end of the data, which favors attackers more than any other location, see Table 3 and [Piet et al., 2023, Yi et al., 2023]. The attack is regarded as successful if and only if the response begins with “Hacked” or “hacked” if not otherwise stated.
@@
We apply SecAlign to five open-weight models [Inc, 2021]: Mistral-7B-Instruct [Jiang et al., 2023], Llama3-8B-Instruct [Dubey et al., 2024], Llama-7B [Touvron et al., 2023b], Mistral-7B [Jiang et al., 2023], Llama3-8B [Dubey et al., 2024]. The first two models have been SFT-ed with their private commercial instruction tuning datasets, so we could directly apply SecAlign on them with their offered delimiters. For Mistral-7B-Instruct, $d_{\text{instruction}}=$”¡s¿[INST] ”, $d_{\text{data}}=$ ” ”, and $d_{\text{response}}=$ ” [/INST]”. For Llama3-8B-Instruct, $d_{\text{instruction}}=$
”¡—begin_of_text—¿¡—start_header_id—¿system¡—end_header_id—¿”,
$d_{\text{data}}=$ ”¡—eot_id—¿¡—start_header_id—¿user¡—end_header_id—¿”, and
$d_{\text{response}}=$ ”¡—eot_id—¿¡—start_header_id—¿assistant¡—end_header_id—¿”. The last three are base pretrained models and should be SFTed before DPO [Rafailov et al., 2024], so we perform standard (non-defensive) SFT following [Chen et al., 2025a], which reserves three special tokens for each of the delimiters. That is, $d_{\text{instruction}}=$[MARK] [INST] [COLN], $d_{\text{data}}=$[MARK] [INPT] [COLN], and $d_{\text{response}}=$[MARK] [RESP] [COLN]. The models have to be used with the exact prompt format, see Section 2.1, that is consistent in our training, otherwise the model performance may drop unpredictably due to the inherent sensitivity to prompt templates in existing LLMs.
@@
In DPO, we use sigmoid activation $\sigma$ and $\beta=0.1$ as the default recommendation. Due to the involvement of two checkpoints $\pi_{\theta},\pi_{\mathrm{ref}}$ in DPO Eq. 3, the memory consumption almost doubles. To ease the training, we adopt LoRA [Hu et al., 2022], a memory efficient fine-tuning technique that only optimizes a very small proportion ($<0.5\%$ in all our studies) of the weights but enjoys performance comparable to fine-tuning the whole model. The LoRA hyperparameters are r=64, lora_alpha=8, lora_dropout=0.1, target_modules = ["q_proj", "v_proj"]. We use the TRL library [von Werra et al., 2020] to implement DPO, and Peft library [Mangrulkar et al., 2022] to implement LoRA. Our training requires 4 NVIDIA Tesla A100s (80GB) to support Pytorch FSDP [Zhao et al., 2023]. We perform DPO for 3 epochs with the tuned learning rates $[1.4,1.6,2.0,1.4,1.6]\times 10^{-4}$ for the five models above respectively. In standard SFT (required before SecAlign for base models) and defensive SFT (the precise StruQ defense [Chen et al., 2025a]), we fine-tune the LLMs for 3 epochs using the learning rate $[20,2.5,2]\times 10^{-6}$ for the three base models above respectively.